Copy of Calculate the pH during the titration of 20.00 mL of 0.1000 M HNO2(aq) with 0.2000 M NaOH(aq) after 7 mL of the base have been added. Ka of nitrous acid = 7.1 x 10-4.
Answer #1

initially

millimoles of HNO2 = 20.00 x 0.100 = 2.0

millimoles of NaOH added = 7.0 x 0.200 = 1.40

after NaOH added

millimoles of HNO2 = 2.0 - 1.40 = 0.60

millimoles of NaNO2 = 1.40

total volume = 20 + 7 = 27.0 mL

[HNO2] = 0.60 / 27.0 = 0.022 M

[NaNO2] = 1.40 / 27.0 = 0.052 M

pH = pKa + log [NaNO2] / [HNO2]

pKa = - log Ka = - log [7.1 x 10-4] = 3.15

pH = 3.15 + log [0.052] / [0.022]

pH = 3.52